OpenStudy (anonymous):

Integrate \[\int \frac{e^x+1}{e^x} dx\]

OpenStudy (anonymous):

that is just \[\int\frac{e^x}{e^x}+\frac{1}{e^x}dx = \int\frac{e^x}{e^x}dx+\int\frac{1}{e^x} =\int1dx + \int\frac{1}{e^x}= x + \int\frac{1}{e^x}dx + C\]

OpenStudy (anonymous):

Ah.. sigh! -.-' I should have seen that. :P

OpenStudy (anonymous):

\[x-e ^{-x} +c\]